The brakes are set on a train running at 30 miles per hour, when ½ miles from a station. The slows down uniformly, coming to rest at the station. Find the acceleration, and the time elapsed in stopping.


1
Expert's answer
2021-03-10T05:53:50-0500

Acceleration =ΔvΔt

                       = 30/60           

                       =0.5 m/s


Time = 0.5 * 0.5

= 0.25

=0.25 * 60

= 15 minutes